OpenStudy (calculusxy):

If points A(2,2), B(4,5), and C(6,2) represents the vertices of a triangle, determine its area. @Mertsj @satellite73 @agent0smith

OpenStudy (anonymous):

Your name is calculusxy. Do you have to use calculus for this or just normal geometry?

OpenStudy (calculusxy):

I put calculusxy because I like to study about calculus. But this is a normal geometry question.

OpenStudy (anonymous):

herons formula would work i think

OpenStudy (anonymous):

Do it intuitively. A = 1/2 * bh b = ((Ax - Bx)^2 + (Ay - By)^2)^(1/2) Let M be equal to the midpoint between A and B, or ((Ax+Bx)/2, (Ay+By)/2). h = ((Mx - Cx)^2 + (My - Cy)^2)^(1/2) A = 1/2 * (((Ax - Bx)^2 + (Ay - By)^2)((Mx - Cx)^2 + (My - Cy)^2))^(1/2)

OpenStudy (calculusxy):

I do not understand this at all :/ Can you just explain it to me in words?

OpenStudy (anonymous):

It's kind of lengthy compared to Heron's formula, you may want to just use that. Or I could explain it.

OpenStudy (anonymous):

do you know heron's formula? it finds the area given the lengths of the sides you can use the distance formula to find the lengths of all three sides, then plug them in to heron's formula i can't write it here, no latex

OpenStudy (calculusxy):

Well, i only know 1/2(b*h)

OpenStudy (anonymous):

Okay, then just use my slightly more intuitive method. We know that the area of a triangle is half the product of the base and height lengths (e.g., A = bh * 1/2). The base length is the distance between A and B. b = ((Ax - Bx)^2 + (Ay - By)^2)^(1/2) or just: sqrt(((Ax - Bx)^2 + (Ay - By)^2)) That is, Ax = 2, Ay = 2, Bx = 4, By = 5. In order to find the height of triangle, we need to find the midpoint of the base. Why? Because an altitude goes from the midpoint to the top, and represents the height of the triangle. If you don't understand, draw this triangle on graph paper and draw a line from the midpoint of the base to the top vertex. Let M be the midpoint. M((Ax+Bx)/2, (Ay+By)/2) Now to find the height, we take the distance between M and C. h = ((Mx - Cx)^2 + (My - Cy)^2)^(1/2) Now substitute b and h into the equation A = bh/2

OpenStudy (calculusxy):

Is Mx and My the midpoints of the x and the y cooridinates?

OpenStudy (anonymous):

Sorry, also make sure the line is orthogonal to the base.

OpenStudy (anonymous):

Yes, M(Mx, My)

OpenStudy (calculusxy):

What is orthogonal?

OpenStudy (anonymous):

Orthogonal = perpendicular.

OpenStudy (anonymous):

Essentially.

OpenStudy (anonymous):

Have you tried to see where these points are? It is a nice triangle The area 3 x 4/2=6

OpenStudy (calculusxy):

Is your way without using the graph paper, because i need such a way?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Just use the formula I gave you.

OpenStudy (calculusxy):

Is there any other easier way to remember the formula that you gave me? I truly found it helpful and would be happy to memorize it easily.

OpenStudy (anonymous):

It is ridiculous. We are making of a very easy problem a very difficult one. That is not Math. Math is to to try to find the easiest way to solve a problem.

OpenStudy (anonymous):

The way my mind works is very programmatic, and it's not a standard formula. It's not as though a geometry teacher has ever stood up and said, "Okay, everybody who wants an A has to memorize this formula." I just derived it in three seconds in my head, then wrote it down. It doesn't take much to derive. You could, alternatively, use Heron's formula and memorize that.

OpenStudy (anonymous):

@eliassab, it is easy. You're just caught up in the symbols. It's actually quite intuitive.

OpenStudy (anonymous):

You do not need it.

OpenStudy (anonymous):

Explain how you would do it, then.

OpenStudy (anonymous):

oops it would certainly help to draw the triangle, then it would be real easy i didn't, my mistake

OpenStudy (anonymous):

unfortunately draw tool is not working

OpenStudy (anonymous):

Enjoy your discussion. I am going for my morning walk in warm South Africa.

OpenStudy (calculusxy):

Wow! Thanks satellite73, you have helped me a lot. you're really great :)

OpenStudy (anonymous):

I said 6 longtime ago

OpenStudy (anonymous):

the picture tells it all use one half base times height base is the distance from (2,2) to (6,2) namely 4 height is pretty clearly 3 one half of two times three is six

OpenStudy (anonymous):

moral of the story... draw the damned picture first!!

OpenStudy (anonymous):

but you aren't always able to draw the picture

OpenStudy (anonymous):

all of these ways are just the same thing in different media